OpenStudy (anonymous):

Which of the following represents 3x - 5y + 10 = 0 written in slope-intercept form? y = -3/5x + 2 y = 3/5x - 2 y = 3/5x + 2

OpenStudy (espex):

\[3x - 5y + 10 = 0 \rightarrow -5y = -3x - 10 \rightarrow y=\frac{-3}{-5}x+\frac{-10}{-5}\rightarrow y=\frac{3}{5}x+2\]

OpenStudy (espex):

Makes sense?

OpenStudy (anonymous):

yes thanks